40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 04:09:53 2011, Koji, Update, LSC, REFL55 PD update REFL55.pdf
    Reply  Mon Jun 20 18:36:49 2011, Suresh, Update, LSC, REFL55 PD update REFL55_response.pdf
Message ID: 4605     Entry time: Tue May 3 04:09:53 2011     Reply to this: 4845
Author: Koji 
Type: Update 
Category: LSC 
Subject: REFL55 PD update 

[Rana, Koji]

REFL55 was modified. The noise level confirmed. The PD is now ready to be installed.

 


Kevin's measurement report told us that something was wrong with REFL55 PD. The transimpedance looked OK, but the noise level was terrible (equivalent to the shotnoise of 14mA DC current).

Rana and I looked at the circuit, and cleaned up the circuit, by removing unnecessary 11MHz notch, 1k shunt resister, and so on.

I made a quick characterization of the PD.

First page:

The transimpedance ws measured as a function of the frequency. The resonance was tuned at 55MHz. The notch was tuned at 110MHz in order to reject the second harmonics. The transimpedance was ~540V/A at 55MHz. (For the calibration, I believed the DC transimpedance of 50V/A and 10000V/A for the DC paths of this PD and #1611, respectively, as well as the RF impedance (700V/A0 of #1611.

Second page:

Output noise levels were measured with various amount of photocurrent using white light from a light bulb. The measurement was perforemed well above the noise level of the measurement instruments.

Third page:

The measured output noise levels were converted into the equivalent current noise on the PD. The dark noise level agrees with the shot noise level of 1.5mA (i.e. 22pA/rtHz). In deed, the noise level went up x~1.5 when the photocurrent is ~1.4mA.

Attachment 1: REFL55.pdf  172 kB  Uploaded Tue May 3 05:10:55 2011  | Hide | Hide all
REFL55.pdf REFL55.pdf REFL55.pdf
ELOG V3.1.3-